OpenStudy (anonymous):

Find the area of a regular hexagon with side length of 5cm. Round your answer to the nearest tenth. 21.7cm^2 32.5cm^2 112.5cm^2 65cm^2

ganeshie8 (ganeshie8):

use the area of hexagon formula :- area of hexagon with side s : A = \(\large \frac{3\sqrt{3}}{2} s^2\)

OpenStudy (anonymous):

\[a={3\sqrt{3}\frac}{ 2 } \]

ganeshie8 (ganeshie8):

? something throwing u off in the formula ha ?

OpenStudy (anonymous):

i can't get the equation box to work

ganeshie8 (ganeshie8):

yea i see that

ganeshie8 (ganeshie8):

just type it i can understand..

OpenStudy (anonymous):

im trying to put A= 3 suqrt 3/2 5^2

ganeshie8 (ganeshie8):

thats correct ! simplify it

ganeshie8 (ganeshie8):

use calculator

OpenStudy (anonymous):

my calculator won't let me put it in

OpenStudy (anonymous):

A=3suqrt 3/2 5^2=25

ganeshie8 (ganeshie8):

use wolframalpha

OpenStudy (anonymous):

D.

ganeshie8 (ganeshie8):

Yes